What is velocity of light in water ?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this Class X Science answer. We will update the answer very soon.
Yash Patari
2 years ago

Speed of light in a water medium (v1) = 2.2×108m/s

Speed of light in glass medium (v2) = 2×108m/s

Refractive index of a medium (n) = Speed of light in water/Speed of light in the glass

n = v1/v2

n = 2.2×108m/s /2×108m/s

n = 1.1